Dear Tutor,
I am having a hard time solving contingency table and answer the following questions. Thank you in advance.
Hair Type Brown Blond Black Red Totals
Wavy 20 5 15 3 43
Straight 80 15 65 12 172
Totals 100 20 80 15 215
Table 3.22
c.What is the probability that a randomly selected child will have either brown or blond hair?
- I am really confused with this question. I would say use P(brown)+P(blond)-P(brown AND blond). 80/215+20/215-(100/120+ 20/120)
d.What is the probability that a randomly selected child will have wavy brown hair?
-43/215
e.What is the probability that a randomly selected child will have red hair, given that he or she has straight hair?
-P(red|straight) = P(red AND straight)/P(straight) = 12/172
f.If B is the event of a child having brown hair, find the probability of the complement of B.
-probability of the complement of B will be all the hair except B. 115/215
g.In words, what does the complement of B represent? Red, blod, black hair.

Answer by stanbon(75887)

Hair Type Brown Blond Black Red Totals
Wavy 20 5 15 3 43
Straight 80 15 65 12 172
Totals 100 20 80 15 215
---------------------
c.What is the probability that a randomly selected child will have either brown or blond hair?
P(brown or blond) = (100+20)/215
-------------------------------------------
d.What is the prob that a randomly selected child will have wavy brown hair?
P(wavy and brown) = 20/215
------------------------------
e.What is the probability that a randomly selected child will have red hair, given that he or she has straight hair?
P(red | straight) = 12/172
------------------------------
f.If B is the event of a child having brown hair, find the probability of the complement of B.
[P(brown)]' = (215-100)/215 = 1-(100/215) = 115/215
------------
g.In words, what does the complement of B represent? Red, blod, black hair.
Depends on what "B" means
-----------
Cheers,
Stan H.
-------------